OpenStudy (unklerhaukus):

Second Order ODE \[\cos^2x\frac{\text d^2y}{\text dx^2}=1\]

OpenStudy (unklerhaukus):

\[\cos^2x\frac{\text d^2y}{\text dx^2}=1\]\[\frac{\text d^2y}{\text dx^2}=\sec^2x\]\[\iint\text d^2y=\iint\sec^2x\cdot\text dx^2\]\[\int y\cdot\text dy =\int\tan x+c\cdot\text dx\]\[\frac 12y^2=\ln \left|\sec x\right|+cx+d\]

OpenStudy (unklerhaukus):

my book has a different answer to this have i missed something?

OpenStudy (turingtest):

what the heck does dx^2 mean o-0 ?

OpenStudy (unklerhaukus):

\[\cos^2x\frac{\text d y \text dy}{\text dx\text dx}=1\]\[\frac{\text d y \text dy}{\text dx\text dx}=\sec^2x\]\[\iint\text dy\cdot\text dy=\iint\sec^2x\cdot\text dx\cdot\text dx\]\[\int y\cdot\text dy =\int\tan x+c\cdot\text dx\]\[\frac 12y^2=\ln \left|\sec x\right|+cx+d\]

OpenStudy (turingtest):

and your answer is right on the RHS, but that 1/2y^2 is wrong

OpenStudy (turingtest):

that is an ill-formed way to write it in my opinion I would just take it one step at a time

OpenStudy (experimentx):

try differentiating it twice ... if you get the DE then it's correct.

OpenStudy (unklerhaukus):

how come the LHS is just \(y\)?

OpenStudy (unklerhaukus):

if there is a simpler way what is it/

OpenStudy (foolaroundmath):

I would write \[\frac{d^{2}y}{dx^{2}} = \frac{d}{dx}(\frac{dy}{dx})\] Then integrate both sides wrt x

OpenStudy (turingtest):

\[\cos^2 xy''=1\]\[\frac{d^2y}{dx^2}=\sec^2x\implies\frac{dy}{dx}=\int\sec^2x dx=\ln|\sec x|+C_1\]seperation of the variables on the second derivative will not give you y, but y' instead

OpenStudy (turingtest):

@FoolAroundMath has stated the formality better that I...

OpenStudy (turingtest):

oops, got ahead of myself above :/\[\cos^2 xy''=1\]\[\frac{d^2y}{dx^2}=\frac{d^{2}y}{dx^{2}} = \frac{d}{dx}(\frac{dy}{dx})=\sec^2x\implies\frac{dy}{dx}=\int\sec^2 x dx=\tan x+C_1\]separate the variables again\[y=\int\tan x+C_1dx=\ln|\sec x|+C_1x+C_2\]

OpenStudy (turingtest):

@FoolAroundMath 's notation should make it more clear why we don't get 1/2y^2 on the LHS

OpenStudy (experimentx):

wolf says correct http://www.wolframalpha.com/input/?i=y%27%27+-+sec^2x+%3D+0

OpenStudy (unklerhaukus):

\[\cos^2x\frac{\text d^2y}{\text dx^2}=1\] \[\frac{\text d^2y}{\text dx^2}=\sec^2x\]\[\frac{\text dy}{\text dx}=\int\sec^2x\cdot\text dx\]\[y =\int\tan x+c\cdot\text dx\]\[=\ln \left|\sec x\right|+cx+d\]

OpenStudy (turingtest):

right, and as I mentioned I think @FoolAroundMath 's notation makes the process even more clear

OpenStudy (unklerhaukus):

a \(\text d\) magically melts away

OpenStudy (turingtest):

no it doesn't, watch closely...

OpenStudy (turingtest):

\[\cos^2x\frac{\text d^2y}{\text dx^2}=1\]\[\frac{\text d^2y}{\text dx^2}=\frac d{dx}\left(\frac{dy}{dx}\right)=\sec^2x\]\[\int d\left(\frac{dy}{dx}\right)=\int\sec^2x dx\]\[\frac{dy}{dx}=\tan x+C_1\]\[\int dy=\int \tan x+C_1dx\]\[y=\ln|\sec x|+C_1x+C_2\]

OpenStudy (turingtest):

notice that by utilizing the notation in the second line it becomes clear that separation of variables is being applied as usual

OpenStudy (unklerhaukus):

I have to wait for this to sink in \[\int \text d\left(\frac{\text dy}{\text dx}\right)=\frac{\text dy }{\text dx}\]

OpenStudy (turingtest):

let\[u=\frac{dy}{dx}\]is not\[\int du=u+C\]?

OpenStudy (turingtest):

we could write this whole thing subbing u for dy/dx

OpenStudy (unklerhaukus):

it make perfect sense with a subsitution \[\int \text du=u+c\]

OpenStudy (unklerhaukus):

yeah it just took a while to get it

OpenStudy (anonymous):

I'm just learning a lot watching this debate ^_^ Perhaps a tutorial on the proper syntax usage for all these \(\frac{d^2x}{dx^2}\) & \(\frac{d}{dx}\) & dx types of things would be in order? (gives @FoolAroundMath a medal for helping out too)

OpenStudy (unklerhaukus):

does\[\iint \text d^2y=y\]

OpenStudy (unklerhaukus):

+c

OpenStudy (turingtest):

yeah, I learned from the user JamesJ (an actual professor), that \[\int\int d^2y\]actually has no meaning it's undefined, and is what is called an ill-formed expression

OpenStudy (unklerhaukus):

oh,

OpenStudy (turingtest):

along with\[\int\int(dx)^2;~~~\int\int dxdx\]and other things like that in short, you do have to be very careful with the notation in cases like this

OpenStudy (unklerhaukus):

why is that ?

OpenStudy (turingtest):

well, what would it mean? nobody has given it one

OpenStudy (turingtest):

if you want to invent a meaning for those expressions above, that's fine, but it won't give you the right answer to your ODE in the way that you have chosen to define it

OpenStudy (unklerhaukus):

it would have worked if \[\iint \text d^2y=y\]

OpenStudy (turingtest):

but you have to analyze what \(d^2y\) means is it \(d(dy)\) ?

OpenStudy (unklerhaukus):

yes

OpenStudy (unklerhaukus):

oh it could be dydy

OpenStudy (unklerhaukus):

i see the confusion

OpenStudy (turingtest):

this is pushing the boundary of my knowledge, but I know for sure that dydy is an ill-formed expression as far as d^2y=d(dy) I'm not sure about that, but if so I guess that would mean that\[\int\int d^2y=\int\int d(dy)=\int dy=y+C\]but then what do you do with \[dx^2\]? is that\[dx^2=d(dx)\]I don't think that works...

OpenStudy (turingtest):

\[\int f(x)dx^2;~~~~~\int\int f(x)dx^2\]don't mean anything to me

OpenStudy (unklerhaukus):

\[\text d^2w=\text d\cdot\text dw\] \[\text dw^2=\text dw\cdot\text dw\]

OpenStudy (experimentx):

$$dx^2 \text{ looks more like } d(x^2) = du \text{ under substitution } u =x^2$$

OpenStudy (turingtest):

@UnkleRhaukus those definitions I suppose make your ODE come out right, so go ahead and use them I guess. I don't know if those are generally accepted by the mathematical community either way, you see how that by defining \(d^2y=d(dy)\) <_(I really think this one should have parentheses to show that d(dy) is the differential of the differential, and no some multiplication) and as I said, I'mm pretty sure that there is some mathematical subtly sin committed by writing\[\int\int f(x)dxdx\]but I am still a student, and would need clarification as to why that is the case

OpenStudy (experimentx):

http://math.stackexchange.com/questions/21199/is-dy-dx-not-a-ratio i guess this would help

OpenStudy (turingtest):

better perhaps if we could find an entry on the possibility of using dxdx for dA in a double integral